How long is a string reaching from the top of a 16- ft pole to a point 13 ft from the base of the pole?
Answer by wilft1(217)

you use the pathagorius theorm for this
a^2+b^2=c^2
so, we have 256+169=c^2
425=c^2
c=20.61 ft